OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10

OpenStudy (anonymous):

x = 2, y = -1

OpenStudy (anonymous):

2x = 1-3y x=1/2-3y/2 Plug this in for x in the second eqn. 3(1/2-3y/2) -4y = 10 Solve for y. Whatever your y-value is, plug this number in to the first eqn to solve for x.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

 Choose the correct classification of by number of terms and by degree.